# Emergence of non-ergodic multifractal quantum states in geometrical fractals

### v1

@@ Introduction
Since the discovery of Anderson’s localization paradigm [1], there has been a significant interest in critical phenomena that transcend the traditional dichotomy of extended and localized states. Among these, a class of novel states, commonly referred to as non-ergodic multifractal states (NEMs) [2, 3, 4], has attracted particular attention. NEMs occupy a finite volume of the system but their support is a set of points that, in the thermodynamic limit, has a zero measure, challenging the conventional ideas of how a quantum state can be distributed in space. The concept of multifractality finds broad applications in a wide range of domains of science and engineering, including fluid dynamics [5], topography [6], seismology [7], finance [8], physiology [9], and epidemic modeling [10]. It also inspired advancements in machine learning, granting unexplored settings to incorporate larger hierarchical depth and achieve stable and robust prediction performances [11, 12]. In condensed matter physics, the concept of NEM is strictly connected to the singular continuous nature of the spectrum [13, 14, 15] and the critical scaling behavior [16, 17, 18] observed at phase transitions, which give rise to anomalous subdiffusive dynamics [19, 20, 21] and make NEM particularly intriguing in the context of quantum transport. Additionally, they exhibit sublinear growth of entanglement entropy [22, 23], defying the standard formulation of eigenstate thermalization hypothesis [24]. These attributes are essential for understanding quantum coherence of eigenstates [25] and many-body localization phenomena [26, 27]. Moreover, it is conjectured that multifractality can enhance electron-electron interactions leading to elevated critical temperature in superconductors [28, 29].
@@
Incommensurate potentials in one-dimensional systems [30, 31, 32, 33] and randomness in higher-dimensional systems [34, 35, 36, 37] are genuine mechanisms to induce disorder causing the emergence of NEMs in quantum systems. However, their direct experimental observations remain problematic due to finite-size effects and the dependence on fine-tuning, posing considerable difficulties for further exploration of NEMs. In this work, we propose an approach to generating such states. Rather than relying on randomness to induce disorder, we introduce a model that generates multifractal states through geometric manipulation by placing substitutional point-like defects in a regular lattice. Specifically, we focus on the case when these defects are arranged to transform the parental lattice into a regular fractal geometry [38]. In the past few years, properties of artificial quantum fractals structures have been systematically explored both experimentally and theoretically [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Here, our aim is to demonstrate how the self-similar structure and non-integer dimensionality of fractals offer a natural framework for designing and studying multifractal eigenstates. We consider single-particle quantum mechanics of a tight-binding model on a Sierpiński Gasket (SG), and through a comprehensive numerical analysis we demonstrate multiple crossovers between extended states and distinct stages of multifractality in the model spectrum. By calculating the fractal dimensions, we identify energy intervals where transitions to multifractal behavior occur. By examining the singularity spectrum, we gain deeper insights into the multifractal nature of individual states, unveiling a variety of behaviors including metallic and insulating NEM. In addition, we observe formation of clusters in the spatial structure of the wavefunctions that we classify as compact localized states (CLSs) [53, 54], which emerge as localized modes on the lattice subsets due to destructive interference on resonant groups of sites.
@@
Recent advancements in spin-resolved scanning tunneling microscopy (STM) and spectroscopy techniques [55, 56, 57] have enabled precise control over quantum states in low-dimensional structures, where quantum confinement is strong. Having this in mind, we also present a detailed numerical analysis of the density-density correlation functions, which can be directly observed in such experiments. By pinpointing the regions where crossovers to NEMs occur, we compute these correlators and show that they exhibit a power-law decay combined with irregular oscillations and revivals in the frequency domain caused by self-similarity of the underlying fractal structure. Overall, these features form a robust criterion to unambiguously distinguish NEMs from both extended and localized states. Yet another verifiable aspect of multifractal systems is the mutual avoidance of eigenstates [58] at large energy separations, offering additional mechanisms by which them stand apart from quantum systems without NEMs in the spectrum.
@@ Results
The degree of delocalization in a system can be probed by examining the structure of its eigenstates in a given basis. A common approach [59, 60, 61] derived from the box-counting method [3, 18] is to analyze the asymptotic scaling properties of moments of order $q$ of the wavefunction $\Phi$ with the number of sites $N$:
$\sum_{i=1}^{N}|\langle{i}|{\Phi}\rangle|^{2q}\sim N^{-D_{q}},$ (2)
@@
The next step is to employ a more refined measure of multifractality – the singularity spectrum $f(\alpha)$, a dual measure of the fractal dimensions (see Methods) with which one can assess the strength of multifractality of a single NEM. In Fig. (3), we present some case studies. The ground state ($E=-3.666$), represented by the red curve, displays a narrow distribution peaked at $\alpha=1$, close to the shape typical for an ideal extended state, as produced, for instance, by a Gaussian Orthogonal Ensemble (thick black line). On the other hand, the singularity spectrum represented with the blue curve for $E=-2.184$ is distributed across a large domain of $\alpha$ values and has its maximum shifted far away from $\alpha=1$, as compared to the extended state. As anticipated already from its vanishingly small fractal dimension $D_{2}$, this picture closely resembles the case of an ideal localized state, with infinite width. The yellow curve depicts the singularity spectrum of the NEM degenerate state at $E=1.847$. Its parabolic shape, as well as the finite width and the apex shifted to $\alpha>1$, are in good agreement with the previous predictions [3].
@@
Using these exemplary cases as a benchmark, one can examine the distribution of the singularity spectrum for different groups of eigenstates and acknowledge the emergence of different degrees of multifractality. NEMs can be not very pronounced, and their singularity spectrum can retain features of metallic or insulating phases [58, 63]. For such “mildy” multifractal states, the shift of the maximum of $f(\alpha)$ may be minor, or its parabolic shape may be distorted, along with other anomalous behaviors known as termination or freezing [3]. One distinct class of states, neither extended nor localized in a rigorous sense, is observed for certain integer energy values. Though these states have a fractal dimension typical for NEMs, their singularity spectra do not abide by the conventional classification. As one can see from Fig. (4), the left branches of the $f(\alpha)$ curves for these states are primarily concentrated around its maximum (like for the metallic states), while the right branches extend in a manner similar to that of the insulating states. This suggests that these states exhibit a form of localization, albeit not restricted to a few sites in a single region, leading us to refer to them as compact localized states (CLS) [64, 54]. The emergence of CLS can be attributed to the interplay of hidden symmetries and destructive interference at resonant sites located along the reflection axes—characteristics inherent to the underlying fractal geometry, as visualized in the amplitude profiles (see the Supplemental Material). In fractal geometries, CLSs were first identified and numerically characterized in [65].
@@
Figure 4: Singularity spectrum for the Sierpiński Gasket: representative cases of compact localized states. In each case, the left side of the $f(\alpha)$ curve clusters around its peak, while the right side extends in a way similar to insulating states. This unique behavior blends features of both extended and localized states, indicating a novel hybrid phase.
@@
We further expand upon our findings by probing the density-density correlators that encode for the overlap between eigenstates and serve as an additional tool to discriminate the NEMs, particularly at critical points when the correlator $\mathcal{F}(\omega,E)$ decays as a power law (see Methods). Fig. (5) illustrates the behavior of $\mathcal{F}(\omega,E)$ for the previously selected representatives of ergodic, localized and NEM states. As predicted by theory, the two-points correlator for the extended (red curve) and localized (blue curve) states exhibits a broad plateau across several orders of $\omega$. The intensity of the correlator reflects the degree of eigenstates overlap: it is low for extended states that share minimal information across a wide spatial region, and high for localized states that are concentrated in a confined region. However, in this particular case of localization, the plateau is especially pronounced, and from the amplitude profile of the state (see the Supplemental Material), it is evident that localization occurs within a single region, though with a slower decay. This suggests that localization is algebraic, rather than of the Anderson type. At large values of energy separation $\omega$, the correlator for both extended and localized states approaches the uncorrelated regime, $\mathcal{F}(\omega,E)=1$, as marked by the typical GOE statistics. The NEM state (yellow curve) qualitatively behaves as expected, displaying a shorter plateau followed by a power-law decay towards the uncorrelated regime as $\mathcal{F}(\omega,E)\sim\left(\omega/E_{Th}\right)^{-\mu}$. The exponent $\mu$ and the Thouless energy $E_{Th}$ are model-dependent and, unlike in disordered quantum systems, the lack of variable parameters in our Hamiltonian makes a possible analytic fit difficult. Notably, the density-density correlator behavior at higher frequencies is marked by pronounced oscillations with clear revivals and dips below the uncorrelated limit. This non-trivial behavior signals the presence of anomalies, reminiscent of effects observed in certain disordered quantum systems stemming from random matrix theory [66, 67, 68]. In our model, these oscillations arise from the competition between the eigenstate avoidance and the resonant states interference driven by the self-similar structure of the Sierpinski Gasket (SG). As shown in [58], the eigenstates avoidance is responsible for dips in $\mathcal{F}(\omega,E)$ falling below the uncorrelated limit. Such dips indicate that, when a state predominantly occupies a site $r$, this site is likely to be vacant for other states: a behavior that is known as eigenstates stratification. On the other hand, revivals are not typical for randomly disordered systems and specific only for the considered case of regular fractal geometry.
@@
Figure 5: Frequency correlator $\mathcal{F}(\omega,E)$ for three characteristic cases: extended (E = -3.666, red), localized (E = -2.184, blue), and multifractal (E = 1.847, yellow). The ideal case of uncorrelated systems as described by the Gaussian Orthogonal Ensemble (GOE) is represented by the dashed black line. A broad plateau at small values of energy separation $\omega$ is a hallmark of both extended and localized states, indicating non-vanishing correlations. In contrast, NEMs exhibit only a narrow plateau, which diminishes to a singular point in the thermodynamic limit, followed by a distinct power-law decay. At larger values of $\omega$, only NEMs display revivals of the correlator, alongside a stratification phenomenon (in the Supplemental Material, examples with more pronounced revivals are presented). According to theory, the transition between phases at small and large energy separations should occur smoothly. However, in our simulations, discontinuities can appear in the correlator for both extended and localized states within the interval $1010^{-1}<\omega<$ 10 due to numerically zero denominators in the computation of the density of states. Lines are guide to eye.
@@ Discussion
With this work, we continue the research line of studying quantum phenomena in artificial fractal geometries, specifically focusing on the case of the Sierpiński gasket. We have shown that in clean fractal, without any added randomness, non-ergodic multifractal states naturally emerge, complementing the prior studies on wavefunction localization [45], quantum transport [21], and eigenstate statistics in fractals, and fractal energy spectra [71]. We performed a multi-tool analysis of both the overall properties of SG spectrum and the specific features of individual NEMs and provided sensitive indicators to distinguish between different types of them. Among other findings, we demonstrated that one particular type of states – compact localized states – respond in a highly symmetric and structured way to introducing random disorder, making them an interesting platform for quantum control.
@@
We suggest that artificial fractal structures – systems that have already been created in a lab – are a natural platform to create and study multifractal states experimentally, which so far – to our knowledge – has been possible only in doped semiconductors [72] and yet remains an ongoing challenge. Our numerical results, particularly on the two-point density-density correlation function $\mathcal{F}(\omega,E)$, provide a direct method to distinguish NEMs using scanning probe microscopy techniques (recent research has also demonstrated applicability of similar methods in acoustic metamaterials [73] and photonic systems [51, 74, 75]).
@@ Methods
Data are available upon reasonable request.
@@ Acknowledgements
We thank C. Morais Smith, M.R. Slot, A. Moustaj, M.Conte and Y. in ’t Veld for the insightful discussions.
This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, grant agreement 854843-FASTCORR. The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@ Introduction
Since the discovery of Anderson’s localization paradigm [1], there has been a significant interest in critical phenomena that transcend the traditional dichotomy of extended and localized states. Among these, a class of novel states, commonly referred to as non-ergodic multifractal states (NEMs) [2, 3], has attracted particular attention. NEMs occupy a finite volume of the system but their support is a set of points that, in the thermodynamic limit, has a zero measure, challenging the conventional ideas of how a quantum state can be distributed in space. The concept of multifractality finds broad applications in a wide range of domains of science and engineering, including fluid dynamics [4], topography [5], finance [6], physiology [7], and epidemic modeling [8]. It also inspired advancements in machine learning, granting unexplored settings to incorporate larger hierarchical depth and achieve stable and robust prediction performances [9, 10]. In condensed matter physics, the concept of NEM is strictly connected to the singular continuous nature of the spectrum [11, 12, 13] and the critical scaling behavior [14] observed at phase transitions, which give rise to anomalous subdiffusive dynamics [15, 16] and make NEM particularly intriguing in the context of quantum transport. Additionally, they exhibit sublinear growth of entanglement entropy [17], defying the standard formulation of eigenstate thermalization hypothesis [18]. These attributes are essential for understanding quantum coherence of eigenstates [19] and many-body localization phenomena [20, 21]. Moreover, it is conjectured that multifractality can enhance electron-electron interactions leading to elevated critical temperature in superconductors [22, 23].
@@
Incommensurate potentials in one-dimensional systems [24, 25] and randomness in higher-dimensional systems [26, 27, 28] are genuine mechanisms to induce disorder causing the emergence of NEMs in quantum systems. However, their direct experimental observations remain problematic due to finite-size effects and the dependence on fine-tuning, posing considerable difficulties for further exploration of NEMs. In this work, we propose an approach to generating such states. Rather than relying on randomness to induce disorder, we introduce a model that generates multifractal states through geometric manipulation by placing substitutional point-like defects in a regular lattice. Specifically, we focus on the case when these defects are arranged to transform the parental lattice into a regular fractal geometry [29]. In the past few years, properties of artificial quantum fractals structures have been systematically explored both experimentally and theoretically [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Here, our aim is to demonstrate how the self-similar structure and non-integer dimensionality of fractals offer a natural framework for designing and studying multifractal eigenstates. We consider single-particle quantum mechanics of a tight-binding model on a Sierpiński Gasket (SG), and through a comprehensive numerical analysis we demonstrate multiple crossovers between extended states and distinct stages of multifractality in the model spectrum. By calculating the fractal dimensions, we identify energy intervals where transitions to multifractal behavior occur. By examining the singularity spectrum, we gain deeper insights into the multifractal nature of individual states, unveiling a variety of behaviors including metallic and insulating NEM. In addition, we observe formation of clusters in the spatial structure of the wavefunctions that we classify as compact localized states (CLSs) [41, 42], which emerge as localized modes on the lattice subsets due to destructive interference on resonant groups of sites.
@@
Recent advancements in spin-resolved scanning tunneling microscopy (STM) and spectroscopy techniques [43, 44] have enabled precise control over quantum states in low-dimensional structures, where quantum confinement is strong. Having this in mind, we also present a detailed numerical analysis of the density-density correlation functions, which can be directly observed in such experiments. By pinpointing the regions where crossovers to NEMs occur, we compute these correlators and show that they exhibit a power-law decay combined with irregular oscillations and revivals in the frequency domain caused by self-similarity of the underlying fractal structure. Overall, these features form a robust criterion to unambiguously distinguish NEMs from both extended and localized states. Yet another verifiable aspect of multifractal systems is the mutual avoidance of eigenstates [45] at large energy separations, offering additional mechanisms by which them stand apart from quantum systems without NEMs in the spectrum.
@@ Results
The degree of delocalization in a system can be probed by examining the structure of its eigenstates in a given basis. A common approach [46, 47] derived from the box-counting method [2, 14] is to analyze the asymptotic scaling properties of moments of order $q$ of the wavefunction $\Phi$ with the number of sites $N$:
$\sum_{i=1}^{N}|\langle{i}|{\Phi}\rangle|^{2q}\sim N^{-(q-1)D_{q}},$ (2)
@@
The next step is to employ a more refined measure of multifractality – the singularity spectrum $f(\alpha)$, a dual measure of the fractal dimensions (see Methods) with which one can assess the strength of multifractality of a single NEM. In Fig. (3), we present some case studies. The ground state ($E=-3.666$), represented by the red curve, displays a narrow distribution peaked at $\alpha=1$, close to the shape typical for an ideal extended state, as produced, for instance, by a Gaussian Orthogonal Ensemble (thick black line). On the other hand, the singularity spectrum represented with the blue curve for $E=-2.184$ is distributed across a large domain of $\alpha$ values and has its maximum shifted far away from $\alpha=1$, as compared to the extended state. As anticipated already from its vanishingly small fractal dimension $D_{2}$, this picture closely resembles the case of an ideal localized state with infinite span of $f(\alpha)$. While the singularity spectrum at $\alpha=0$ does not reach zero here, as it should for strictly localized states in the thermodynamic limit, scaling analysis of the fifth and sixth generations of the gasket (see Supplemental Material) suggest a convergence trend. While finite-size scaling corrections likely affect these results, this evaluation is consistent with complementary indicators, including the two-point correlation function and the amplitude profiles, all of which support the interpretation of this state as localized. Finally, the yellow curve depicts the singularity spectrum of one of the NEM degenerate states at $E=1.847$. Its parabolic shape, as well as the finite width and the apex shifted to $\alpha>1$ remain consistent under size scaling as shown in the Supplemental Material, aligning well with the previous predictions [2].
@@
Using these exemplary cases as a benchmark, one can examine the distribution of the singularity spectrum for different groups of eigenstates and acknowledge the emergence of different degrees of multifractality. NEMs can be not very pronounced, and their singularity spectrum can retain features of metallic or insulating phases [45, 49]. For such “mildy” multifractal states, the shift of the maximum of $f(\alpha)$ may be minor, or its parabolic shape may be distorted, along with other anomalous behaviors known as termination or freezing [2]. One distinct class of states, neither extended nor localized in a rigorous sense, is observed for certain integer energy values. Though these states have a fractal dimension typical for NEMs, their singularity spectra do not abide by the conventional classification. As one can see from Fig. (4), the left branches of the $f(\alpha)$ curves for these states are primarily concentrated around its maximum (like for the metallic states), while the right branches extend in a manner similar to that of the insulating states. This suggests that these states exhibit a form of localization, albeit not restricted to a few sites in a single region. Despite coexistence of several regions of localization, we still refer to them as compact localized states (CLS) [41]. The emergence of CLS can be attributed to the interplay of hidden symmetries and destructive interference at resonant sites located along the reflection axes—characteristics inherent to the underlying fractal geometry, as visualized in the amplitude profiles (see the Supplemental Material). In fractal geometries, CLSs were first identified and numerically characterized in [50].
@@
Figure 4: Singularity spectrum for the Sierpiński Gasket: representative cases of of compact localized states. While a CLS with a single cluster of non-zero $|\psi|^{2}$ would exhibit behavior akin to a standard localized state, the coexistence of several clusters causes the left side of the $f(\alpha)$ curve to concentrate densely around the GOE peak, with its density reflecting the number of isolated regions within the CLS. Further analysis of such states can be found in Supplemental Material.
@@
We further expand upon our findings by probing the density-density correlators that encode for the overlap between eigenstates and serve as an additional tool to discriminate the NEMs, particularly at critical points when the correlator $\mathcal{F}(\omega,E)$ decays as a power law (see Methods). Fig. (5) illustrates the behavior of $\mathcal{F}(\omega,E)$ for the previously selected representatives of ergodic, localized and NEM states. As predicted by theory, the two-points correlator for the extended (red curve) and localized (blue curve) states exhibits a broad plateau across several orders of $\omega$. The intensity of the correlator reflects the degree of eigenstates overlap: it is low for extended states that share minimal information across a wide spatial region, and high for localized states that are concentrated in a confined region. However, in this particular case of localization, the plateau is especially pronounced, and from the amplitude profile of the state (see the Supplemental Material), it is evident that localization occurs within a single region, though with a slower decay. This suggests that localization is algebraic, rather than of the Anderson type. At large values of energy separation $\omega$, the correlator for both extended and localized states approaches the uncorrelated regime, $\mathcal{F}(\omega,E)=1$, as marked by the typical GOE statistics. The NEM state (yellow curve) qualitatively behaves as expected, displaying a shorter plateau followed by a power-law decay towards the uncorrelated regime as $\mathcal{F}(\omega,E)\sim\left(\omega/E_{Th}\right)^{-\mu}$. The exponent $\mu$ and the Thouless energy $E_{Th}$ are model-dependent and, unlike in disordered quantum systems, the lack of variable parameters in our Hamiltonian makes a possible analytic fit difficult. Notably, the density-density correlator behavior at higher frequencies is marked by pronounced oscillations with clear revivals and dips below the uncorrelated limit. This non-trivial behavior signals the presence of anomalies, reminiscent of effects observed in certain disordered quantum systems stemming from random matrix theory [52, 53]. In our model, these oscillations arise from the competition between the eigenstate avoidance and the resonant states interference driven by the self-similar structure of the Sierpinski Gasket (SG). As shown in [45], the eigenstates avoidance is responsible for dips in $\mathcal{F}(\omega,E)$ falling below the uncorrelated limit. Such dips indicate that, when a state predominantly occupies a site $r$, this site is likely to be vacant for other states: a behavior that is known as eigenstates stratification. On the other hand, revivals are not typical for randomly disordered systems and specific only for the considered case of regular fractal geometry.
@@
Figure 5: Frequency correlator $\mathcal{F}(\omega,E)$ for three characteristic cases: extended (E = -3.666, red), localized (E = -2.184, blue), and multifractal (E = 1.847, yellow). The ideal case of uncorrelated systems as described by the Gaussian Orthogonal Ensemble (GOE) is represented by the dashed black line. A broad plateau at small values of energy separation $\omega$ is a hallmark of both extended and localized states, indicating non-vanishing correlations. In contrast, NEMs exhibit only a narrow plateau, which diminishes to a singular point in the thermodynamic limit, followed by a power-law decay highlighted with the black crosses. At larger values of $\omega$, only NEMs display revivals of the correlator, alongside a stratification phenomenon (in the Supplemental Material, examples with more pronounced revivals are presented). According to theory, the transition between phases at small and large energy separations should occur smoothly. However, in our simulations, discontinuities can appear in the correlator for both extended and localized states within the interval $1010^{-1}<\omega<$ 10 due to numerically zero denominators in the computation of the density of states. Lines are guide to eye.
@@ Discussion
With this work, we continue the research line of studying quantum phenomena in artificial fractal geometries, specifically focusing on the case of the Sierpiński gasket. We have shown that in clean fractal, without any added randomness, non-ergodic multifractal states naturally emerge, complementing the prior studies on wavefunction localization [35], quantum transport [16], and eigenstate statistics in fractals, and fractal energy spectra [56, 32]. We performed a multi-tool analysis of both the overall properties of SG spectrum and the specific features of individual NEMs and provided sensitive indicators to distinguish between different types of them. Among other findings, we demonstrated that one particular type of states – compact localized states – respond in a highly symmetric and structured way to introducing random disorder, making them an interesting platform for quantum control.
@@
We suggest that artificial fractal structures – systems that have already been created in a lab – are a natural platform to create and study multifractal states experimentally, which so far – to our knowledge – has been possible only in doped semiconductors [57] and yet remains an ongoing challenge. Our numerical results, particularly on the two-point density-density correlation function $\mathcal{F}(\omega,E)$, provide a direct method to distinguish NEMs using scanning probe microscopy techniques (recent research has also demonstrated applicability of similar methods in acoustic metamaterials [58] and photonic systems [40, 59]).
@@ Methods
Data are available from the corresponding author upon reasonable request.
@@ Acknowledgements
We are grateful to C. Morais Smith, M.R. Slot, A. Moustaj, M.Conte and Y. in ’t Veld for insightful discussions. We would also like to thank I.M. Khaymovich for valuable correspondence. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, grant agreement 854843-FASTCORR.